OpenStudy (anonymous):

solve the equation. radical 3x-5=4 is the correct answer 3 my options are A.7 B.3 C.1/3 D.49

jimthompson5910 (jim_thompson5910):

sqrt(3x - 5) = 4 3x - 5 = 4^2 3x - 5 = 16 3x - 5+5 = 16+5 3x = 21 3x/3 = 21/3 x = 7 So the answer is choice A

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

what about when i have a equation like this: radical 6m-5=m

jimthompson5910 (jim_thompson5910):

sqrt(6m - 5) = m 6m - 5 = m^2 0 = m^2 - 6m + 5 m^2 - 6m + 5 = 0 Now use the quadratic formula to solve for m m = (-b+-sqrt(b^2-4ac))/(2a) m = (-(-6)+-sqrt((-6)^2-4(1)(5)))/(2(1)) m = (6+-sqrt(36-(20)))/(2) m = (6+-sqrt(16))/2 m = (6+sqrt(16))/2 or m = (6-sqrt(16))/2 m = (6+4)/2 or m = (6-4)/2 m = 10/2 or m = 2/2 m = 5 or m = 1 So the possible solutions are m = 5 or m = 1 You would need to check if they are really indeed solutions. I'll let you do this.

OpenStudy (anonymous):

my options are A. 1 B.5 C. 1,5 or D no real solution

jimthompson5910 (jim_thompson5910):

choice C, when you do check the solutions, you'll find that both 1 and 5 are solutions

OpenStudy (anonymous):

ok thank you

jimthompson5910 (jim_thompson5910):

np